Lyndsay Delello, Fabiana Martinez, Megan Norton – Cultural Cognition and How It Relates to Climate Change.

Your presentation to the class was mostly good, although there were still some coordination issues (overlap of content) between Fabiana’s part and Lyndsay’s part. The fact that you had so many slides also created some time pressure for Meg’s part, so you ended up being rushed.

The major strength of your project was the way you were able to design the survey, get it approved, arrange to have it administered to the residents of Mulledy via Moodle, and do some analysis on the results. It was great that you got started early on this. I think the data from the responses are also quite interesting. If it’s OK with you, I would like to hold onto the Excel spreadsheet and possibly use this in other ways in future years.

The paper is clearly three separate sections spliced together rather than a single text. I know you had a lot of issues to deal with, though, what with Lyndsay’s absence during the time the project was “going” and other things. So I’m not penalizing the group for that.

But there are also some technical issues. The major thing there is that you have not included in the paper a complete bibliography of the sources you consulted. I hope that you didn’t misunderstand the terms of the assignment. Even though I asked you to do the annotated bibliography earlier, you were still supposed to include a complete bibliography in the final paper since, as was true for several groups, additional sources might have emerged as work on the project proceeded. Plus, citations in the text should be checkable from the paper itself (without referring to what you submitted earlier). I can guess how that probably happened, but that just is not acceptable for a research paper like this. (Be aware, too, that in a lot of courses here at HC, problems with citations would carry extremely steep penalties on the overall grade for the paper.)

Specific Comments

1) It would have been good to start the paper with an overall introduction rather than jumping into the first section.

2) I’m not sure what “projects a continuation” in the second sentence means. Do you mean that “it is projected that this will continue?”

3) Maybe “natural processes” here. The idea is that more or less natural processes like digestion in large animals also add methane to the atmosphere. (I say “more or less natural” because the characteristics of these animals and their numbers have been strongly influenced by humans.)

4) The discussion of the reasons for skepticism is not very clear. Do you mean that some people have used the fact that scientific theories are always subject to revision if different evidence is found in later experiments to sow doubt about things (as in “evolution is only a theory, not established fact?”)

5) We saw some specific examples of what you said about US temperature trends in the temperature data that Mike and Drew Bechard analyzed.

6) It’s important to realize that the relation between higher average temperatures and higher precipitation is not uniform, though. Global climate change is affecting different areas in different ways. Some parts of the world (e.g. northern Africa) have actually be getting drier.

6.5) You say small effects can build up over time. That is certainly true, but there’s also the point that even a 1-degree change in global average temperature can have significant effects.

7) The Red Scare may have been an over-reaction, but there’s also no denying that many communist regimes have been extremely repressive and brutal to their populations.

8) Technical point about citations: When referring to an article with multiple authors (like (The Tragedy of the Risk Perception Commons), it would be common to use the abbreviation Kahan et al. to refer to the authors. (This is from Latin words meaning literally “and the rest.”)

Final Project Grades

Presentation: B+ (88)

Paper Content: A- (92)

Paper Form: B (85)